For a data set of the pulse rates for a sample of adult females, the lowest pulse rate is 35 beats per minute, the mean of the listed pulse rates is x=79

minute, and their standard deviation is s = 23.9 beats per minute.

a. What is the difference between the pulse rate of 35 beats per minute and the mean pulse rate of the females?

b. How many standard deviations is that [the difference found in part (a)?

c. Convert the pulse rate of 35 beats per minutes to a z score.

d. If we consider pulse rates that convert to z scores between - 2 and 2 to be neither significantly low nor significantly high, is the pulse rate of 35 beats

significant?


1
Expert's answer
2022-01-14T04:53:37-0500

a) Let a be a lowest pulse rate, then a = 35. In that case the sought value is x - a = 79 - 35 = 44


b) This is "{\frac {44} {23.9}}=1.841" standard deviations


c) z-score for a = 35 is "{\frac {a-x} s}={\frac {35-79} {23.9}}=-1.841"


d) Obtained z-score for 35 pulse rate is lying inside interval (-2;2), which means it is not considered as significantly low